OpenStudy (anonymous):

x,y≥0 f(xf(y))f(y)=f(x+y) f(2)=0 f(x)≠0 0≤x<2 find f(1)

OpenStudy (anonymous):

Is this function only defined for x between 0 and 2?

OpenStudy (anonymous):

no [0,inf]

OpenStudy (anonymous):

look at the first condition

OpenStudy (anonymous):

Ah, alright.

OpenStudy (anonymous):

I've gotten that f(0) = 1 by letting both x = y = 0, but not much beyond that.

OpenStudy (anonymous):

yeap but i am looking for f(1) ok go ahead.

OpenStudy (anonymous):

I see that f(f(1) = 0, so f(1) could be 2, but it could be other things as well.

OpenStudy (kinggeorge):

I'm going to go out on a limb, and say that \(f(1)=2\).

OpenStudy (anonymous):

f(f(1))*

OpenStudy (kinggeorge):

If we let x=y=1, then we have that \(f(1\cdot f(1))=f(2)\). Thus, \(1\cdot f(1)=2\). It follows that \(f(1)=2\).

OpenStudy (kinggeorge):

I did that wrong. one moment.

OpenStudy (anonymous):

yeah, I just got that too. f(x) = 0 for all x greater than 2.

OpenStudy (anonymous):

0 no cause of x=1 y=2 f(3)=0 you can see also f(x)=0 for all x>=2

OpenStudy (kinggeorge):

We have \(f(1\cdot f(1))\cdot f(1)=f(2)=0\) So either \(f(1\cdot f(1))=0\) or \(f(1)=0\). Since we're told \(f(x)\neq 0\), it must be that \(f(1\cdot f(1))=0\)

OpenStudy (kinggeorge):

\(f(x)\neq 0\) for \(0\leq x <2\)

OpenStudy (anonymous):

ok then f(1)>=2

OpenStudy (anonymous):

what is exactly?

OpenStudy (anonymous):

If you let x = 1 and y = f(1), then you get:\[f(1\cdot f(f(1))f(f(1))=f(1+f(1))\Longrightarrow f(f(f(1)))f(f(1))=f(1+f(1))\]Since f(f(1))=0, we have:\[f(f(1)+1)=0\Longrightarrow f(1)+1\ge 2\Longrightarrow 1\le f(1)\le 2\]getting close >.<

OpenStudy (anonymous):

No wait, now i did something wrong >.< one sec.

OpenStudy (anonymous):

it just has to be bigger than 1, not less that 2, scratch that.

OpenStudy (anonymous):

only >=1 we knew that!

OpenStudy (anonymous):

>.< haha

OpenStudy (anonymous):

certainly you have wonderful insight!

OpenStudy (anonymous):

Just trying random things, but nothing is creating a solution >.<

OpenStudy (kinggeorge):

Do you know if there's supposed to be a specific number?

OpenStudy (anonymous):

i think somebody who immediately understand wrong way is successful.dont you?

OpenStudy (anonymous):

no there is not.

OpenStudy (anonymous):

some help! try with x=2/f(1)

OpenStudy (kinggeorge):

I'm sorry, but I've got nothing else on this problem, and I've got to go. Good luck.

OpenStudy (anonymous):

good luck

OpenStudy (anonymous):

x,y≥0 f(xf(y))f(y)=f(x+y) f(2)=0 f(x)≠0 0≤x<2 find f(1) f(f(1))f(1)=f(2) f(f(1))=0 f(1)>=2 or f(1)<0 f(0f(1))f(1)=f(1) f(0)f(1)=f(1) f(0)=1 Great I am not going anywhere..... Let me try more